If 5.43 mol of an ideal gas has a pressure of 6.09 atm and a volume of 40.95 L, what is the temperature of the sample?
Answer #1

Given Values are

n = number of moles of gas present in sample = 5.43 mol

P = Pressure of gas = 6.09 atm

V =Volume it occupies = 40.95 L

T = ?

We know the ideal gas law equation

PV = nRT

where

R = 0.0821 L.atm/mol.K

so by using given values and above formula , lets calculate for temperature

PV = nRT

then

T = PV / nR

= 6.09 atm \times 40.95 L / (5.43 mol \times 0.0821 L.atm/mol.K)

= 559.40 K

So our answer is 559.40 K